Sandy and James work at a local restaurant. Sandy can fold 30 napkins in 10 minutes, while James can fold 39 napkins in 15 minutes. How long will it take them to fold 112 napkins altogether?
Answer by checkley71(8403)

SANDY FOLDS 30 IN 10 MINUTES THEN HE FOLDS 3 PER MINUTE
JAMES FOLDS 39 IN 15 MINUTES THEN HE FOLDS 39/15=2.6 PER MINUTE
TOGETHER THEY FOLD 3+2.6=5.6 PER MINUTE
THUS TO FOLD 112 NAPKINS IT WILL TAKE THEM
112/5.6=20 MINUTES.
PROOF
20*3+20*2.6=60+52=112